OpenStudy (anonymous):

Find the first derivative for y * sqrt(x) + x * sqrt(y) = 9

OpenStudy (anonymous):

\[y'\sqrt{x}+\frac{1}{2\sqrt{x}}y+\sqrt{y}+\frac{xy'}{2\sqrt{y}}=0\] is a start

OpenStudy (anonymous):

now you get to solve that mess for \[y'\]

OpenStudy (anonymous):

I don't understand where the \[(xy')\div2\sqrt{y}\] came from.

myininaya (myininaya):

\[[y \sqrt{x}]'=(y)'\sqrt{x}+y(\sqrt{x})'=y'\sqrt{x}+y \cdot \frac{1}{ 2 \sqrt{x}}\] \[[x \sqrt{y}]'=(x)'\sqrt{y}+x(\sqrt{y})'=1 \sqrt{y}+x \frac{1}{2 \sqrt{y}} \cdot y'\]

myininaya (myininaya):

\[(\sqrt{y})'=(y^\frac{1}{2})'=\frac{1}{2}y^{\frac{1}{2}-1}y'=\frac{1}{2}y^\frac{-1}{2}y'=\frac{1}{2} \frac{1}{y^\frac{1}{2}}y' \]

myininaya (myininaya):

don't forget chain rule is derivative of outside times derivative of inside

OpenStudy (anonymous):

Oh okay I get why.

OpenStudy (anonymous):

Thanks.

myininaya (myininaya):

np :)